OpenStudy (anonymous):

5/7 . 1/5 + 2/7 . 5/4

Directrix (directrix):

5/7 x 1/5 + 2/7 x 5/4 = 5/35 + 10/28 = 1/7 + 5/14 = 2/14 + 5/14 = 7/14 = 1/2.